A binomial event has n = 60 trials. The probability of success on each trial is 0.4. Let X be the count of successes of the event during the 60 trials. What are mean and variance of X?

A) 24, 3.79
B) 24, 14.4
C) 4.90, 3.79
D) 4.90, 14.4
E) 2.4, 3.79
Answer #1

Here, n = 60 , p = 0.4

mean = np
= 60 * 0.4 = 24

varaince = npq
= 60 * 0.4 * 0.6 = 14.4


Option b) is answer